OpenStudy (anonymous):

the arcsin (- root 3/2) Find the exact value of the expression. (Enter your answers in radians.)

OpenStudy (anonymous):

ask yourself if you know a number (angle) between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose sine is \(-\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

240 and 300

OpenStudy (anonymous):

if that is not obvious look at the unit circle on the last page of the attached cheat sheet, and find the angles corresponding to the point where the second coordinate is \(-\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

it is neither of those if you are working in degrees, you have to pick an angle between \(-90\) and \(90\)

OpenStudy (anonymous):

4pi/3 ND 5PI/3

OpenStudy (anonymous):

ok lets go slow first of all it is only one number, not two it is a well defined function, so there cannot be two answers, only one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 4pi/3

OpenStudy (anonymous):

no you have to look on the right side of the unit circle for \(\sin^{-1}(x)\)

OpenStudy (anonymous):

then 5pi/3?

OpenStudy (anonymous):

now i see on the cheat sheet that the angle given on the right side is \(\frac{5\pi}{3}\) but that is not right either it must be between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\)

OpenStudy (anonymous):

so think of a negative angle that is coterminal with \(\frac{5\pi}{3}\)

OpenStudy (anonymous):

the correct answer is \(-\frac{\pi}{3}\)

OpenStudy (anonymous):

that is the only angle between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose sine is \(-\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

if you are working in degrees, it would be \(-60\)